OpenStudy (anonymous):

if f(x)=1/2x-2, what is f(14)?

OpenStudy (anonymous):

If it is f(14) = 1/2x -2 you would have to solve for x, therefore the answer should be 32 . (:

OpenStudy (math&ing001):

Actually all you have to do is plug in 14 like this : f(14) = 1/2 * 14 - 2

OpenStudy (mkmkasim):

1/2(14) = 7 7-2=5 5 is the answer

OpenStudy (mkmkasim):

do you get it?

OpenStudy (anonymous):

yes thank you! :)

OpenStudy (mkmkasim):

no problem!